### v3

# DualSpeechLM: Towards Unified Speech Understanding and Generation via Dual Speech Token Modeling with Large Language Models

Abstract
Extending pre-trained text Large Language Models (LLMs)’s speech understanding or generation abilities by introducing various effective speech tokens has attracted great attention in the speech research community. However, building a unified speech understanding and generation model still faces the following challenges: (1) Due to the huge modality gap between speech and text tokens, extending text LLMs to unified speech LLMs relies on large-scale paired data for fine-tuning, and (2) Generation and understanding tasks prefer information at different levels, e.g., generation benefits from detailed acoustic features, while understanding favors high-level semantics. This divergence leads to difficult performance optimization in one unified model. To solve these challenges, in this paper, we present two key insights in speech tokenization and speech language modeling. Specifically, we first propose an Understanding-driven Speech Tokenizer (USTokenizer), which extracts high-level semantic information essential for accomplishing understanding tasks using text LLMs. In this way, USToken enjoys better modality commonality with text, which reduces the difficulty of modality alignment in adapting text LLMs to speech LLMs. Secondly, we present DualSpeechLM, a dual-token modeling framework that concurrently models USToken as input and acoustic token as output within a unified, end-to-end framework, seamlessly integrating speech understanding and generation capabilities. Furthermore, we propose a novel semantic supervision loss and a Chain-of-Condition (CoC) strategy to stabilize model training and enhance speech generation performance. Experimental results demonstrate that our proposed approach effectively fosters a complementary relationship between understanding and generation tasks, highlighting the promising strategy of mutually enhancing both tasks in one unified model. [^1] Code and demo: https://github.com/lavendery/UUG.
Introduction
Recent advancements in autoregressive large language models (LLMs) have demonstrated excellent performance in the natural language processing community (achiam2023gpt; touvron2023llama; team2023gemini). Leveraging the powerful foundations of text LLMs, recent advancements have led to emergence of speech LLMs that possess speech understanding and generation capabilities. In addition to fine-tuning textual LLMs to separately perform speech understanding tasks (gong2023joint; tang2023salmonn; chu2023qwenaudio; wang2024blsp; wang2025unisep) and generation tasks (wang2023neural; anastassiou2024seed; kim2024clam; wang2024speechx; yang2024uniaudio; yang2025simplespeech; jia2025ditar), developing unified models that excel at both capabilities has been explored in recent years (zhang2023speechgpt; xie2024miniomni2; fu2024vita; nguyen2025spirit; defossez2024moshi; xu2025qwen2). However, several limitations still remain.
First, adapting pre-trained text LLMs to unified speech LLMs still relies heavily on large-scale paired speech-text data (zhang2023speechgpt; defossez2024moshi; xie2024miniomni2; xu2025qwen2). For example, SpeechGPT (zhang2023speechgpt) and SpiritLM (nguyen2025spirit) require approximately 70K and 570K hours of paired data, respectively. This dependence stems from the substantial modality gap between speech and text, which hinders capability transfer. Second, existing speech LLMs struggle to meet the distinct informational needs of understanding and generation. As shown in Figure 1 (a) Left and (b) Left, the Baseline model—trained solely with acoustic tokens—exhibits a contradiction between tasks, i.e., improving one often leads to degradation of the other, highlighting its inability to balance both tasks effectively. In fact, generation tasks demand rich acoustic details (e.g., prosody, emotion, speaker traits) for high-fidelity synthesis (zeghidour2021soundstream; defossez2022high; kumar2023high; wang2023neural), which acoustic tokens capture well but lack high-level semantics (shi2024espnet; dhawan2024codec; anastassiou2024seed; 10447929). Conversely, understanding tasks benefit from semantic features (borsos2023audiolm; rubenstein2023audiopalm; maiti2024voxtlm), but semantic tokens inevitably compromise the acoustic details needed for natural speech generation.
To solve these problems, we propose two key insights from the perspectives of speech tokenization and language modeling. First, we present an Understanding-driven Speech Tokenizer (USTokenizer) that can extract high-level semantic features, which are critical for understanding tasks. Unlike prior methods that rely on only self-supervised learning (SSL) representation quantization (hsu2021hubert; chen2022wavlm) or automatic speech recognition (ASR)-based objectives (du2024cosyvoice; zeng2024glm4) to capture semantics, our approach directly aligns speech tokenizer with the semantic understanding capabilities of text LLMs. This leads to USTokens that have inherently better alignment with text modality, significantly easing modality alignment when adapting text LLMs to speech LLMs. Secondly, building on USTokenizer, we introduce DualSpeechLM, a novel dual speech token modeling framework that effectively handles the distinct informational needs of understanding and generation within a unified end-to-end architecture. Unlike conventional approaches that use the same token as input and output of LLM, our model separates them by using USTokens as input and acoustic tokens as output. Specifically, the USTokenizer provides high-level semantic information to boost understanding tasks, while an AcousticGPT module restores fine-grained acoustic details, enabling diverse and realistic speech generation within a unified end-to-end framework. In addition, we introduce a semantic supervision loss and a Chain of Condition (CoC) strategy to stabilize training and further improve the performance of the unified framework. In summary, our contributions are as follows:
  • We present a USTokenizer that can extract high-level semantic information and reduce the modality gap when adapting text LLMs to speech LLMs.
  • We propose an end-to-end dual token modeling framework, DualSpeechLM, that simultaneously accepts USTokens as input and generates acoustic tokens as output, effectively accommodating distinct informational needs.
  • We also propose a novel semantic supervision loss and a CoC strategy to improve training stability and generation performance.
  • Experiments demonstrate that our method achieves faster convergence and excellent performance with small-scale data, and enhances mutual improvements between understanding and generation tasks.
(a) Generation performance ($\downarrow$), measured by generated speech WER (%), using different amounts of generation (left) or understanding (right) data, with the amount of understanding (left) or generation (right) data fixed as 300K samples.
(b) Understanding performance ($\downarrow$), measured by speech recognition WER (%), using different amounts of generation (left) or understanding (right) data, with the amount of understanding (left) or generation (right) data fixed as 300K samples.
Figure 1: Comparison of baseline and our model on generation and understanding tasks with different ratios of generation and understanding training data.
Related Work
Speech Tokenization
The success of autoregressive language models (achiam2023gpt; touvron2023llama; team2023gemini) has spurred progress in speech LLMs (chu2023qwenaudio; wang2024speechx; wang2024freeze; defossez2024moshi), where speech tokenizers are essential for converting continuous signals into discrete tokens (yangalmtokenizer). Speech tokenizers are typically categorized as acoustic or semantic (borsos2023audiolm; parkerscaling; yanguniaudio15). Acoustic tokens, optimized for signal reconstruction (zeghidour2021soundstream; defossez2022high; yang2023hifi; wang2023neural), capture detailed acoustic features beneficial for generation, but perform poorly on understanding tasks like ASR (shi2024espnet; dhawan2024codec; anastassiou2024seed; 10447929). Previous semantic tokenizers were trained in two ways: (1) applying clustering (borsos2023audiolm; zhang2023speechgpt; shi2023multi) or vector quantization (VQ) (huang2024repcodec) to representations of SSL models  (hsu2021hubert; chen2022wavlm), and (2) applying a VQ layer to the intermediate layer of ASR models  (du2024cosyvoice; zeng2024glm4; du2024cosyvoice2). Although these semantic tokenizers have shown benefits for understanding tasks (borsos2023audiolm; rubenstein2023audiopalm; maiti2024voxtlm), they do not explicitly consider the alignment to the modality of text LLM (li2025baichuanaudio). Furthermore, multi-codebook designs (zhang2023speechtokenizer; defossez2024moshi; yangalmtokenizer) aim to capture semantics and acoustics in different codebooks jointly, but often introduce complexity when integrated with LLMs. In this work, we present a single VQ-codebook USTokenizer, which not only incorporates high-level semantic information but also achieves better alignment between speech and text modality when applied to LLMs.
Speech Language Models
Recent advances in speech LLMs have explored unified understanding and generation (zhang2023speechgpt; pan2024auto; defossez2024moshi; nguyen2025spirit). Some speech LLMs (yang2024uniaudio; shi2025balancing) adopt acoustic tokens to ensure high-fidelity speech synthesis. However, such tokens usually lead to degraded performance in understanding tasks, particularly in low-resource scenarios. By contrast, semantic tokens perform better in understanding tasks. Yet, their lack of acoustic details often results in reduced generation quality. To compensate for generation, existing works (polyak2021speech; zhang2023speechgpt; du2024cosyvoice; nguyen2025spirit) introduce additional components such as diffusion models (ho2020denoising) or flow matching (lipman2022flow), to convert speech tokens into a Mel spectrogram, and then a HiFi-GAN (kong2020hifi-gan) vocoder is used to synthesize waveform with the Mel spectrogram as input. These multi-stage pipelines increase complexity and risk error accumulation (jia2025ditar). To address these limitations, we propose DualSpeechLM, an end-to-end dual-token modeling framework that explicitly models USTokens as input for understanding and acoustic tokens as output for generation, effectively achieving distinct informational requirements.
Methodologies
We propose a novel unified speech LLM framework, comprising an understanding-driven speech tokenization method and a dual-token modeling paradigm. This framework enhances both understanding and generation capabilities in the resulting DualSpeechLLM. In the following, we describe each module of our framework.
USTokenizer
Figure 2: The architecture of USTokenizer, which can extract high-level semantic features aligned with text LLMs via an understanding-driven loss.
As shown in Figure 2, our Understanding-driven Speech Tokenizer (USTokenizer) consists of a pre-trained Whisper encoder, another downsampling Encoder, a vector quantizer (VQ), an upsampling Decoder, an Adapter module, and a frozen text LLM. Given a speech utterance $\mathbf{X}$, we use the final hidden states $\mathbf{E}$ from the Whisper-medium encoder  [^2] https://github.com/openai/whisper as the input to the downsampling Encoder and then the VQ to obtain the Understanding-driven Speech Token (USToken). Additionally, we novelly project the quantized vector from VQ to the input space of LLM via an adapter to integrate the guidance provided by text LLMs to optimize the speech token during training, as described in the following sections.
Encoder and Decoder
The Encoder converts Whisper features $\mathbf{E}$ into representations $\mathbf{H}$ using a 2× downsampling convolution followed by two residual convolutional blocks, then discretized by VQ. Symmetrically, the Decoder reconstructs $\mathbf{\hat{E}}$ from quantized vectors $\mathbf{\hat{H}}$. Both Encoder and Decoder adopt similar residual convolutional blocks with RepCodec (huang2024repcodec). We then compute the reconstruction loss via mean squared error (MSE) between $\mathbf{E}$ and $\mathbf{\hat{E}}$, formulated as:
$\begin{aligned} \mathcal{L}_{\text{reconstruction}}(\mathbf{E},\mathbf{\hat{E}})=\frac{1}{N}\sum_{i=1}^{N}(E_{i}-\hat{E_{i}})^{2}\end{aligned},$ (1)
where $N$ is the number of elements in the embedding vector.
Figure 3: DualSpeechLM’s dual-token modeling paradigm. The left illustrates the baseline pipeline treating LLM input/output as identical tokens. In contrast, our DualSpeechLM (right) incorporates an Acoustic GPT module into the text LLM module for joint training, separately processing USToken inputs and acoustic token outputs through distinct modeling paths, effectively capturing the different levels of information required for both generation and understanding tasks.
Model LLM Input Token Training Data(hrs) ASR S2TT SER SQA
Clean Other En2De - -
$w\downarrow$ $b4\uparrow$ $acc\uparrow$ $b4\uparrow/gs\uparrow$
SpeechGPT  (zhang2023speechgpt) LLaMA-7B D(Hubert) Full FT 70K 42.73 78.54 1.07 - 3.58/40
SpiritLM  (nguyen2025spirit) LLaMA-7B D(Hubert) Full FT 570K 6.0 11.0 - - -
Mini-Omni2  (xie2024miniomni2) Qwen2-0.5B C(Whisper) Full FT 9K 4.8 9.8 - - -
VITA  (fu2024vita) Mixtral-8x7B C(Mel) Full FT 20K 8.14 18.41 - - 7.62/70
Moshi  (defossez2024moshi) Helium-7B D(Mimicodec) Full FT 7M 5.7 - - - -
Qwen2.5-Omni  (xu2025qwen2) Qwen2.5-7B C(Whisper) Full FT $\star$ 1.8 3.4 30.2 60.03 4.28/60
Baseline-Acoustic Phi3.5-3B D(Wavtokenizer) LoRA 4.5K 36.52 80.06 1.91 54.90 17.68/76
Baseline-Semantic Phi3.5-3B D(Hubert) LoRA 4.5K 5.70 14.32 11.13 51.91 42.01/85
DualSpeechLM Phi3.5-3B D(Hubert) LoRA 4.5K 5.56 14.62 10.20 51.77 42.59/85
D(USToken) LoRA 4.5K 4.22 9.71 19.74 60.92 44.38/88
Table 1: Understanding capability evaluation. ‘D’: Discrete, ‘C’: Continuous, ‘Full FT’: full fine-tuning. $w$, $b4$, gsgs denote WER, BLEU4, ChatGPTScore. Qwen2.5-Omni uses 300B audio tokens, and DualSpeechLM uses 405M speech tokens.
Model LLM Input Token Training TTS VC T2ST SC
Clean Other VCTK Es2En Fr2En -
$s\uparrow/w\downarrow/d\uparrow$ $s\uparrow/w\downarrow/d\uparrow$ $b4\uparrow$ $b4\uparrow/gs\uparrow$
GT - - 1.0/3.94/3.86 1.0/5.35/3.78 1.0/2.64/3.62 - - -/-
SpeechGPT LLaMA-7B D(Hubert) Full FT -/22.15/3.97 -/24.55/3.96 - 14.62 14.74 3.50/54
Mini-Omni2 Qwen2-0.5B C(Whisper) Full FT - - - - - 2.22/54
Moshi Helium-7B D(Mimicodec) Full FT - - - - - 1.75/50
Qwen2.5-Omni Qwen2.5-7B C(Whisper) Full FT -/3.73/4.10 -/4.29/4.08 - - - 6.70/70
Baseline-Acoustic Phi3.5-3B D(Wavtokenizer) LoRA 0.88/22.11/3.76 0.87/26.38/3.69 0.80/22.07/3.30 8.52 7.62 0.52/62
Baseline-Semantic Phi3.5-3B D(Hubert) LoRA 0.80/21.72/3.29 0.81/22.32/3.26 0.81/18.88/3.25 18.05 15.76 10.44/60
DualSpeechLM Phi3.5-3B D(Hubert) LoRA 0.81/20.80/3.35 0.81/17.12/3.38 0.82/18.70/3.37 17.54 15.97 11.06/65
D(USToken) LoRA 0.90/9.25/3.86 0.88/9.88/3.82 0.80/10.16/3.46 26.77 23.82 16.24/67
Table 2: Generation capability evaluation. Data usage is the same as Data(hrs) in Table 1. $s$, $w$, $d$, $b4$, gsgs denote SIM, WER, DNSMOS, BLEU4, ChatGPTScore. Qwen2.5-Omni lacks standalone TTS inference, so the Qwen-TTS API is used instead.
Model TTS ($Q\uparrow/S\uparrow$) VC ($Q\uparrow/S\uparrow$)
GT $20_{\pm 0.33}/3.94_{\pm 0.42} 4.4$$48_{\pm 0.23}/4.60_{\pm 0.25}$
SpeechGPT $3.53_{\pm 0.51}$/- -
Qwen-TTS $4.49_{\pm 0.21}$/- -
Baseline-Acoustic $3.67_{\pm 0.48}$$\mathbf{3.77_{\pm 0.52}}$ $3.63_{\pm 0.33}$$3.43_{\pm 0.42}$
Baseline-Semantic $2.26_{\pm 0.53}$$2.76_{\pm 0.45}$ $2.75_{\pm 0.0.56}$$2.93_{\pm 0.57}$
Ours-Hubert $3.48_{\pm 0.61}$$3.59_{\pm 0.47}$ $3.54_{\pm 0.34}$$3.18_{\pm 0.36}$
Ours-USToken $\mathbf{3.89_{\pm 0.31}}$$3.74_{\pm 0.24}$ $\mathbf{3.80_{\pm 0.28}}$$\mathbf{3.55_{\pm 0.37}}$
Table 3: Subjective evaluation on generation tasks. $Q,S$ denote QMOS, SMOS. Ours-Hubert and Ours-USToken refer to DualSpeechLM using Hubert and USToken as input tokens, respectively. Configurations are the same as Table 2.
Vector Quantization
The VQ module discretizes continuous feature representations $\mathbf{H}$ by mapping them to a codebook of learned vectors. In this work, we use a single VQ layer to quantize the feature $\mathbf{H}$ into discrete tokens $\mathbf{Z}$. Following previous works (huang2024repcodec), we add a commitment loss $\mathcal{L}_{\text{commit}}$ to ensure training stability, with more details provided in Appendix B.
Understanding-Driven Loss
The LLM module aims to align the speech token $\mathbf{Z}$ with the LLM’s input space. By optimizing the understanding tasks upon the text LLM, the required understanding capability will be backpropagated to the optimization of speech token, such that the modality gap between speech and text tokens can be effectively reduced. The LLM-based understanding loss is formulated as the likelihood of generating the target response given a speech prompt using a text LLM (e.g., generating answers given the prompt of speech question):
$\mathcal{L}_{\text{Under}}=-\sum_{t=1}^{L}\log p(S_{t}\mid\mathbf{m},S_{<t};\theta),$ (2)
where $L$ is the length of the sequence, $S_{t}$ represents the target token at position $t$, $S_{<t}$ is the sub-sequence of text tokens before $t$, and $\mathbf{m}$ is the feature of speech prompt extracted by USTokenizer. $\theta$ represents parameters of the text LLM.
The LLM module aligns the VQ space with text LLM’s input space, which means USTokens are mapped to an LLM-compatible feature space, enhancing subsequent unified understanding and generation modeling. The final total loss of USTokenizer is as follows:
$\mathcal{L}_{\text{USTokenizer}}=\alpha\cdot\mathcal{L}_{\text{commit}}+\beta\cdot\mathcal{L}_{\text{Under}}+\gamma\cdot\mathcal{L}_{\text{reconstruction}},$
where $\alpha,\beta$ and $\gamma$ are weighting hyperparameter.
DualSpeechLM
As shown in Figure 3, unlike prior Speech LLMs that use the same token for both input and output (yang2024uniaudio; wang2024speechx; shi2025balancing), DualSpeechLM introduces a novel dual-token design by modeling USTokens as input and acoustic tokens as output via an integrated AcousticGPT, effectively accommodating different levels of information required for understanding and generation tasks.
For speech understanding tasks, the USTokenizer first encodes raw speech into USTokens, which are combined with task-specific prompts and fed into the text LLM. The model is trained using Cross-Entropy (CE) loss between predicted and ground-truth text tokens. During inference, the model generates text tokens conditioned on USTokens and prompts. The text tokens are then decoded into the final outputs.
For the generation task, let $\mathbf{U^{tar}}$ and $\mathbf{A^{tar}}$ denote USTokens and acoustic tokens of target speech, $\mathbf{U^{in}}$ and $\mathbf{P}$ represent the USTokens and prompt of input speech. As shown in Figure 3, the text LLM first predicts $\mathbf{U^{tar}}$ conditioned on $\mathbf{P}$ and $\mathbf{U^{in}}$, formulated as:
$p(\mathbf{U^{tar}}|\mathbf{P},\mathbf{U^{in}};\theta)=\prod_{t=1}^{L}p(\mathbf{U^{tar}}_{t}|\mathbf{U^{tar}}_{<t},\mathbf{P},\mathbf{U^{in}};\theta),$ (3)
where $L$ is the length of $\mathbf{U^{tar}}$, $\mathbf{U^{tar}}_{t}$ is the target USToken at position $t$ and $\mathbf{U^{tar}}_{<t}$ represents all previous USTokens. $\theta$ denotes model parameters. To improve training stability and model performance, we introduce a semantic supervision loss:
$\mathcal{L}_{\text{semantic}}=-\frac{1}{L}\sum_{t=1}^{L}\log p(\mathbf{U^{tar}}_{t}|\mathbf{U^{tar}}_{<t},\mathbf{P},\mathbf{U^{in}};\theta)$ (4)
Next, the predicted USToken sequence $\mathbf{U^{tar}}$ is passed to AcousticGPT, which autoregressively generates acoustic tokens. They are then converted into speech via a Speech Decoder. For both understanding and generation, we adopt text LLM as backbone and fine-tune it using parameter-efficient Low-Rank Adaptation (LoRA) (hu2022lora). Within the unified framework, we integrate AcousticGPT into text LLM to enable joint end-to-end training. This allows DualSpeechLM to simultaneously capture high-level semantics for understanding and fine-grained acoustic details for generation, which we refer to as dual-token modeling. Notably, the acoustic tokens and Speech Decoder are modular and can be implemented using any open-source codec model.
AcousticGPT
Inspired by AudioLM (borsos2023audiolm), we design an AcousticGPT module to autoregressively generate acoustic tokens conditioned on semantic hidden states and speaker embeddings. Built on a GPT-style transformer, the model consists of six causal blocks, each containing a causal self-attention layer followed by a Multi-Layer Perceptron (MLP). The multi-head self-attention captures dependencies under causal constraints. In addition, each block includes residual connections and layer normalization to enhance training stability and model performance. Further details are provided in Appendix B.
Chain of Condition
To improve training robustness, we introduce a Chain of Condition (CoC) strategy for AcousticGPT. Specifically, given the dynamically changing semantic hidden states, we randomly sample conditioning signals from three sources with equal probability: prompt hidden states $\mathbf{S}_{p}$, predicted USTokens $\mathbf{U^{tar}}$, or their concatenation $\mathbf{S}_{p};\mathbf{U^{tar}}]$. During inference, we consistently use the concatenated form $\mathbf{S}_{p};\mathbf{U^{tar}}]$. This CoC strategy acts as a regularizer, preventing overfitting to any single modality and improving the model’s adaptability by exposing it to varied conditioning signals during training. Importantly, it also mitigates the adverse effects of inaccurate USToken predictions by allowing the model to flexibly rely on more stable prompt hidden states or concatenation. As a result, CoC improves the alignment between USTokens and text tokens in the shared latent space, fostering stronger synergy between understanding and generation representations. The acoustic loss in AcousticGPT is defined as:
$\mathcal{L}_{\text{acoustic}}=-\sum_{t=1}^{L}\log p(\mathbf{A^{tar}}_{t}|\mathbf{A^{tar}}_{<t},\mathbf{S},\mathbf{Spk};\theta),$ (5)
where $\mathbf{S}\in\{\mathbf{S}_{p},\mathbf{U^{tar}},[\mathbf{S}_{p};\mathbf{U^{tar}}]\}$ is selected via CoC, $\mathbf{Spk}$ is the speaker embedding, and $\mathbf{A^{tar}}_{t}$ is the predicted acoustic token at time $t$. $p(\mathbf{A^{tar}}_{t}|\mathbf{A^{tar}}_{<t},\mathbf{S},\mathbf{Spk};\theta)$ is the predicted probability of generating $\mathbf{A^{tar}}_{t}$, given $\{\mathbf{A^{tar}}_{<t},\mathbf{S},\mathbf{Spk}\}$. Finally, the overall generation loss combines semantic and acoustic objectives as:
$\mathcal{L}_{\text{generation}}=\lambda\cdot\mathcal{L}_{\text{semantic}}+\xi\cdot\mathcal{L}_{\text{acoustic}},$ (6)
where $\lambda,\xi$ are weighting hyperparameters.
Experiment Setup
Dataset
We train the USTokenizer using multiple speech understanding tasks, including Automatic Speech Recognition (ASR), Speech Emotion Recognition (SER), and Speech Question Answering (SQA). They are based on LibriSpeech (7178964), IEMOCAP (busso2008iemocap) and SQA dataset from SALMONN (tang2023salmonn), respectively. In the SQA dataset, both questions and answers were automatically generated based on LibriSpeech transcripts using ChatGPT, forming spoken-question–text-answer pairs.
DualSpeechLM is trained on eight speech understanding and generation tasks using approximately 4,500 hours of speech data. For understanding, we use LibriSpeech for ASR task, the English-to-German (En2De) subset of CoVoST2 (wang2020covost) for Speech-to-Text Translation (S2TT) task, IEMOCAP (busso2008iemocap) for the SER task, and the SQA dataset (tang2023salmonn) constructed from LibriSpeech using ChatGPT. For generation, we train Text-to-Speech (TTS) on LibriTTS-R (koizumi2023librittsr), Text-to-Speech Translation (T2ST) on Spanish-to-English (Es–En) and French-to-English (Fr–En) subsets of CVSS (jia2022cvss), and Voice Conversion (VC) on a 1,000-hour subset of LibriHeavy (kang2023libriheavy) with evaluation on 400 pairs from VCTK (Yamagishi2019CSTRVC). Additionally, the Speech Conversation (SC) task is built by synthesizing speech from the SQA dataset using the Volcengine TTS API. Please see Appendix C for dataset details.
Evaluation Metrics
Objective Evaluation
For evaluation of the understanding tasks, we use Word Error Rate (WER) for ASR, BLEU-4 for S2TT, accuracy (ACC) for SER, and both BLEU-4 and ChatGPTScore for SQA, respectively. For the generation tasks, we employ speaker similarity (SIM), WER, and DNSMOS to assess TTS and VC performance. BLEU-4 is used for T2ST. Both BLEU-4 and ChatGPTScore are used for evaluating the SC task.
Subjective Evaluation
We also conduct subjective evaluations for generation tasks, focusing primarily on TTS and VC. For each task, subjective assessment is carried out from speech quality (QMOS) and speaker similarity (SMOS). More details can be found in Appendix C.
Model Settings
In experiments, we compare four variants under the same configurations: (1) Baseline-Acoustic, where both input and output are acoustic token from WavTokenizer (ji2024wavtokenizer); (2) Baseline-Semantic, where both input and output are HuBERT token (hsu2021hubert); (3) DualSpeechLM-Hubert, using HuBERT token as input, acoustic token from WavTokenizer as output; and (4) DualSpeechLM-USToken, using our USToken as input, acoustic token from WavTokenizer as output. Notably, the Baseline-Acoustic model requires 160k steps to converge under multi-task settings, while others converge in 60k steps. Further details are provided in the Appendix D.
Results and Analyses
Understanding and Generation Performance
Model LLM Token ASR S2TT SER SQA
Clean Other En2De - -
$wer\downarrow$ $bleu4\uparrow$ $acc\uparrow$ $bleu4\uparrow$
Ours Phi3.5-3B Hubert 5.56 14.62 10.20 51.77 42.59
USToken 4.22 9.71 19.74 60.92 44.38
Ours Vicuna-7B Hubert 5.28 16.64 10.92 54.08 42.68
USToken 4.15 9.69 20.46 58.42 44.60
Table 4: The understanding performance comparison of DualSpeechLM when using different text LLM backbones.
Model LLM Input Token TTS VC T2ST SC
Clean Other VCTK Es2En Fr2En -
$s\uparrow/w\downarrow/d\uparrow$ $s\uparrow/w\downarrow/d\uparrow$ $\uparrow bleu4bleu4\uparrow$
DualSpeechLM Phi3.5-3B Hubert 0.81/20.80/3.35 0.81/17.12/3.38 0.82/18.70/3.37 17.54 15.97 11.06
USToken 0.90/9.25/3.86 0.88/9.88/3.82 0.80/10.16/3.46 26.77 23.82 16.24
DualSpeechLM Vicuna-7B Hubert 0.87/13.58/3.88 0.86/15.56/3.80 0.78/19.86/3.38 22.29 19.74 13.17
USToken 0.89/12.63/3.90 0.87/14.65/3.86 0.78/17.03/3.43 29.58 25.40 17.18
Table 5: Comparison of generation performance of DualSpeechLM using different text LLM backbones.
Model TTS VC T2ST SC
Clean Other VCTK Es2En Fr2En -
$s\uparrow/w\downarrow/d\uparrow$ $s\uparrow/w\downarrow/d\uparrow$ $\uparrow bleu4bleu4\uparrow$
DualSpeechLM 0.90/9.25/3.86 0.88/9.88/3.82 0.80/10.16/3.46 26.77 23.82 16.24
USTokenizer Ablation
w/o Understanding-driven Loss 0.90/8.59/3.86 0.88/9.45/3.81 0.82/9.90/3.39 24.03 22.73 15.59
w/o Reconstruction Loss 0.83/52.50/3.02 0.81/54.24/2.96 0.79/26.08/3.33 32.43 29.46 17.09
DualSpeechLM Ablation
w/o Semantic Loss 0.80/167.56/3.85 0.85/175.11/3.83 0.80/264.53/3.45 0.09 0.07 0.15
w/o CoC 0.89/9.96/3.84 0.87/10.34/3.81 0.79/13.06/3.38 - - -
Table 6: Ablation study using generation tasks. We use Phi-3.5-3B as the text LLM backbone.
Model ASR S2TT SER SQA
Clean Other En2De - -
$wer\downarrow$ $bleu4\uparrow$ $acc\uparrow$ $bleu4\uparrow$
DualSpeechLM 4.22 9.71 19.74 60.92 44.38
USTokenizer Ablation
w/o Understanding-driven Loss 4.81 10.43 14.81 52.7 37.67
w/o Reconstruction Loss 4.74 10.46 18.5 45.21 46.44
DualSpeechLM Ablation
w/o Semantic Loss 4.31 9.61 18.67 60.35 43.86
Table 7: Ablation study using understanding tasks. Phi-3.5-3B is used as the text LLM backbone.
We first compare the performance of our method with prior work. The results for understanding and generation are shown in Table 1, 2 and 3, respectively. From these results, we observe the following:
(1) USToken has better modality commonality with text, reducing the difficulty of modality alignment when adapting text LLMs to speech LLMs. DualSpeechLM-USToken outperforms both Baselines and DualSpeechLM-Hubert across almost all tasks, highlighting USToken’s enhanced semantic understanding capabilities. Additionally, on translation tasks (S2TT and T2ST), which require the inherent translation capabilities of text LLM, our DualSpeechLM-USToken still significantly outperforms both Baselines and DualSpeechLM-Hubert. This demonstrates that USToken has better alignment with text modality, allowing it to retain more capabilities of text LLM during training. Overall, both DualSpeechLM-Hubert and DualSpeechLM-USToken exhibit better performance than the baseline, indicating that the design of Acoustic GPT in DualSpeechLM also helps alleviate the pressure of text LLM. The comparison of convergence speed is provided in Appendix A.
(2) DualSpeechLM achieves strong performance in both tasks with small-scale resources. Specifically, compared to Baselines, both DualSpeechLM-Hubert and DualSpeechLM-USToken demonstrate significant performance improvements, highlighting the effectiveness of DualSpeechLM. Compared to existing unified models, which rely on larger datasets and full fine-tuning, such as SpeechGPT and Mini-Omni2, DualSpeechLM-USToken achieves better performance with just 4.5k hours of data using parameter-efficient fine-tuning, LoRA.
(3) DualSpeechLM simultaneously fulfill the distinct information requirements of generation and understanding by dual-token modeling. Comparisons between Baseline-Acoustic and Baseline-Semantic in Table 1, 2, and 3 show that semantic tokens excel in understanding tasks, while acoustic tokens perform better in generation, e.g., higher SIM and DNSMOS scores in TTS. However, acoustic tokens carry weaker semantics reflected by higher WER results and lower translation quality. In contrast, our DualSpeechLM demonstrates superior performance on both understanding and generation tasks simultaneously due to its dual-token modeling. Another evidence is shown in Figure 1, where our method consistently improves performance on both generation and understanding tasks as the amount of either type of training data increases. This demonstrates its strong ability to support one task without compromising the other.
Performance with Different LLM Backbones
To further assess the generalizability of DualSpeechLM and the effectiveness of USToken across different backbones, we compare models using various text LLMs as backbones. Table 4 and 5 show that DualSpeechLM-USToken consistently outperforms DualSpeechLM-Hubert across almost all tasks, regardless of the underlying LLMs. This demonstrates that USToken not only aligns more effectively with text LLMs but also transfers well across model architectures. Additionally, our DualSpeechLM achieves performance comparable to or better than existing methods, with different text LLMs as the backbone. This highlights the strong compatibility and robustness of DualSpeechLM across different LLM backbones, further demonstrating its effectiveness as a unified framework for both speech understanding and generation.
Ablation Study
We perform ablation studies to evaluate the contribution of each component, with results shown in Table 6 and 7.
The Influence of Understanding-driven Loss.
We ablate the understanding-driven loss in USTokenizer while keeping all other components unchanged. As shown in Table 7, removing this LLM-based objective leads to significant performance drops in understanding tasks (e.g., BLEU4 decreases from 44.38 to 37.67 on SQA), highlighting its importance for aligning USTokens more closely with the semantic space of text LLMs using the understanding-driven loss. For T2ST and SC in Table 6, we also observe a performance drop after removing the understanding-driven loss. This can be attributed to these two tasks rely more heavily on the internal capabilities of the text LLM. The understanding-driven loss plays a critical role in enhancing modality commonality between USTokens and text tokens, thereby helping to preserve as much of the original ability of the text LLM as possible when adapting the text LLM to speech-related tasks.
Effect of the Reconstruction Loss.
The third row in Table 6 and 7 shows the results when the reconstruction loss is removed from USTokenizer. For tasks like T2ST that rely more on the reasoning capabilities of the text LLM, performance actually improves. This suggests that removing the reconstruction objective pushes USTokenizer to better align with the text modality, thus narrowing the modality gap and allowing the LLM to operate more effectively on semantically rich tasks. However, we observe notable performance degradation in SER, TTS, and VC tasks, indicating that the reconstruction loss is essential for preserving fine-grained information, which is crucial for these tasks.
Effect of Semantic Loss.
The fourth row in Table 6 and 7 shows the results of removing semantic loss from DualSpeechLM. This leads to a slight decline in understanding performance, indicating that reconstructing USTokens, as enforced by semantic loss, provides useful semantic supervision that benefits understanding tasks. More critically, we observe a substantial performance drop across all generation tasks. Without semantic loss, the text LLM struggles with producing high-quality USTokens, resulting in degraded inputs to AcousticGPT and ultimately poor generation of acoustic tokens. These results underscore the pivotal role of semantic loss in ensuring accurate semantic representation, which is essential not only for understanding but also for maintaining high-fidelity generation in unified DualSpeechLM.
Influence of CoC.
The final row in Table 6 shows the impact of removing the Chain of Condition (CoC) strategy from the AcousticGPT module, which is applied only to TTS and VC tasks. Performance drops notably on both tasks, indicating that CoC’s stochastic conditioning improves alignment between DualTokens and text tokens in the shared latent space. This enhanced alignment leads to more stable and accurate acoustic token generation.
Conclusions
In this work, we aim to develop a unified speech LLM that can simultaneously excel in both speech understanding and generation. We first propose USTokenizer, which reduces the modality gap between speech and text when adapting text LLMs to speech LLMs. Based on this, we present an end-to-end DualSpeechLM that effectively accommodates the different informational requirements for understanding and generation by a dual-token modeling strategy. Experiments indicate that our method achieves significantly better performance than baselines, enabling mutual improvements between understanding and generation. In the future, we plan to improve DualSpeechLM to larger, more diverse multilingual and cross-domain datasets to further explore its generalization and robustness.
Appendix A Appendix
Appendix B A. Convergence Speed Comparison
Figure 4 shows the training loss curves of adapting text LLMs to speech LLMs with different tokens as input. The speech LLMs with USTokenizer and HuBERT converge faster and achieve lower final loss than other Speech LLMs using WavTokenizer. Notably, USTokenizer achieves the fastest convergence at the lowest loss, underscoring its effectiveness in adapting text LLMs to speech LLMs. This advantage stems from USTokenizer’s alignment with the semantic space of text LLMs, enabling USTokens to share stronger modality commonality with text and thus facilitating more efficient learning.
Figure 4: Training loss curves for the understanding task. From top to bottom, the curves correspond to models trained with WavTokenizer, HuBERT, and USTokenizer as input, respectively.
Appendix C B. Model Details
Vector Quantization
In this work, we use a single quantizer. As shown in the architecture of USTokenizer in our main paper, the input $\mathbf{H}$ is first reshaped and passed through the codebook, where the distance between the input and the codebook vectors is computed. The nearest codebook vector is selected for each input element, producing the quantized output $\mathbf{Z}$. The commitment loss in VQ modules encourages the encoder’s continuous outputs to stay close to the codebook vectors, stabilizing training. Based on this, we define the commitment loss as:
$\mathcal{L}_{\text{commit}}=\left\|\mathbf{H}-\text{sg}(\mathbf{Z})\right\|_{2}^{2},$ (7)
where $\mathbf{H}$ is the continuous feature vector from the encoder, $\mathbf{Z=vq(H)}$ is the quantized vector from the codebook, $\text{sg}(\cdot)$ represents stop-gradient operator, that treats $\mathbf{Z}$ as a constant during backpropagation, and $\|\cdot\|_{2}$ is the L2-norm. By this commitment loss, the encoder parameters are updated to move $\mathbf{H}$ closer to $\mathbf{Z}$ during training.
AcousticGPT
Figure 5: The details of AcousticGPT .
The AcousticGPT module is implemented as a GPT-style structure, which autoregressively predicts next acoustic tokens based on previous acoustic tokens and conditioned on semantic representations and speaker embeddings. As shown in Figure 5, the module consists of six stacked causal blocks, each comprising a causal multi-head self-attention layer followed by a multi-layer perceptron (MLP). Within each block, layer normalization is applied before both the self-attention and MLP modules, and residual connections are added after each to facilitate gradient flow and improve training stability. The MLP component includes a linear layer, GELU activation, another linear projection, and dropout for regularization. A final layer normalization layer is applied after the last causal block. During training, the hidden states extracted from text LLMs are first selected using the CoC strategy and then concatenated with speaker embeddings extracted by a pre-trained 3D-Speaker encoder (chen20243d) to form the input to the AcousticGPT module. Then the model generates acoustic tokens that are coherent in both content and speaker characteristics. This architecture enables AcousticGPT to produce high-fidelity speech representations aligned with semantic and stylistic information.
Appendix D C. DataSet Statistics
Training Data Statistics
Task Description Data Source #Hours #Samples
Understanding
ASR Automatic Speech Recognition Librispeech (7178964) 960 280K
S2TT Speech-to-Text Translation CoVoST2-En2De (wang2020covost) 430 290K
[…]
SQA Speech Question Answering SQA (tang2023salmonn) 960 280K
Generation
TTS Text-to-Speech LibriTTS-R (koizumi2023librittsr) 960 350K
T2ST Text-to-Speech Translation CVSS (jia2022cvss)-Es2En 60 70K
CVSS (jia2022cvss)-Fr2En 109 130K
VC Voice Conversion Libriheavy (kang2023libriheavy) 1K 290K
SC Speech Conversation Synthetic SQA (tang2023salmonn) 100 28K
Total 4.5K 1.7M
Table 8: Training data statistics.
Task Description Data Source Eval Metrics
Understanding
ASR Automatic Speech Recognition Librispeech test clean WER
Librispeech test other
S2TT Speech-to-Text Translation CoVoST2-En2De BLEU4
SER Speech Emotion Recognition IEMOCAP Session 5 ACC
SQA Speech Question Answering SQA (tang2023salmonn) BLEU4/ChatGPTScore
Generation
TTS Text-to-Speech LibriTTS-R test clean SIM/WER/DNSMOS
LibriTTS-R test other
T2ST Text-to-Speech Translation CVSS-Es2En BLEU4
CVSS-Fr2En
VC Voice Conversion VCTK (Yamagishi2019CSTRVC) SIM/WER/DNSMOS
SC Speech Conversation Synthetic SQA (tang2023salmonn) BLEU4/ChatGPTScore
Table 9: Test data statistics for DualSpeechLM.
During the training of USTokenizer, we utilize a dataset comprising nearly 2,000 hours of speech, jointly optimizing for multiple spoken language understanding tasks, including Automatic Speech Recognition (ASR), Speech Emotion Recognition (SER), and Speech Question Answering (SQA). Specifically, the LibriSpeech (7178964) train set was used for ASR, while the IEMOCAP (busso2008iemocap) Sessions 1–4 are employed for SER. For the SQA task, we utilize a dataset from SALMONN (tang2023salmonn), where questions and answers are automatically generated from the LibriSpeech transcripts using ChatGPT. In this setting, the LLM is required to generate text responses to prompts by jointly considering both the speech input and the corresponding text question. We refer to this task as Speech Question Answering (SQA).
As shown in Table 8, in DualSpeechLM, we jointly train a unified model for both speech understanding and generation across eight distinct tasks, leveraging approximately 4,500 hours of speech data. For speech understanding, we use the LibriSpeech (7178964) dataset to train and evaluate the Automatic Speech Recognition (ASR) task; the English-to-German (En2De) split of CoVoST2 (wang2020covost) for training and evaluating the Speech-to-Text Translation (S2TT) task; IEMOCAP (busso2008iemocap) Sessions 1–4 for training and Session 5 for evaluating the Speech Emotion Recognition (SER) task; Similarly, in the SQA task, our DualSpeechLLM is required to generate text responses to prompts by jointly considering both the speech input and the corresponding text question. We use the SQA dataset (tang2023salmonn), constructed from LibriSpeech using ChatGPT, for both training and evaluation. Specifically, 400 data pairs are randomly selected from the dataset to form a separate test set, which does not overlap with the training set. For speech generation, we employ the LibriTTS-R (koizumi2023librittsr) dataset for training and evaluating the Text-to-Speech (TTS) task; the Spanish-to-English (Es–En) and French-to-English (Fr–En) splits of  (jia2022cvss) for training and evaluating the Text-to-Speech Translation (T2ST) task. For the Voice Conversion (VC) task, we randomly select 1,000 hours from LibriHeavy (kang2023libriheavy) for training. To evaluate the VC task, we randomly construct 400 data pairs from VCTK (Yamagishi2019CSTRVC) by pairing utterances of the same sentence spoken by different speakers, as well as different utterances by the same speaker for speaker embeddings. Furthermore, we randomly sample 100 hours of speech from the SQA training set and use the open-source Volcengine TTS API to synthesize both corresponding questions and answers into speech, forming speech-question–speech-answer pairs for the Speech Conversation (SC) task. For evaluation, we construct a separate SC test set by randomly selecting 400 additional data pairs and applying the same synthesis procedure. In the SC task, the DualSpeechLLM is required to generate speech responses to speech prompts that include speech input and the corresponding speech question.
Objective Evaluation Metrics
USTokenizer Configuration
Model Configuration
Codebook size 1024
Codebook dimension 1024
VQ layers 1
Encoder dimension 1024
Decoder dimension 1024
Optimization Configuration
Global batch size 32
Optimizer AdamW
Optimizer hyperparameter $\beta_{1}=0.9,\beta_{2}=0.999$
Warmup Steps 3000
Peak learning rate 3e-5
Minimum learning rate 1e-5
Learning rate decay 5e-2
Numerical precision bf16
Table 10: USTokenizer Configuration.
As shown in Table 9, we adopt a set of objective metrics to evaluate both understanding and generation performance in DualSpeechLM. Specifically, Word Error Rate (WER) is used in ASR, TTS, and VC to quantify transcription errors by calculating the proportion of substitutions, insertions, and deletions relative to the total number of words in the reference, where lower WER indicates higher transcription fidelity. For TTS and VC tasks, we first transcribe the generated speech using Whisper large-v3 (radford2023robustwhisper)  [^3] https://huggingface.co/openai/whisper-large-v3 and then compute WER based on the transcribed output. BLEU4 (papineni2002bleu) is employed in S2TT, T2ST, SQA, and SC to measure the n-gram overlap (up to 4-grams) between generated and reference texts, where higher scores reflect stronger textual alignment, especially useful for translation. For SQA and SC, since the answers are typically generated based on the given speech input, BLEU4 can be used to evaluate the consistency between the generated and reference answers. However, as BLEU4 has limited coverage for diverse but correct expressions, we additionally adopt ChatGPTScore for further evaluation for SQA and SC tasks. Specifically, we use ChatGPT 4.1 (achiam2023gpt) to directly rate the relevance, fluency, and correctness of generated and reference answers, with the evaluation prompt detailed as shown in Figure 6. For the AC task, we first use Whisper large-v3 (radford2023robustwhisper) to transcribe speech answers into text, and then employ the same evaluation prompt with ChatGPT 4.1 to assess the results. For the SER task, we use classification accuracy (ACC) to assess the proportion of correctly predicted emotion categories. For speaker similarity (SIM), used in TTS and VC, we employ the wavlm-base-plus-sv (chen2022wavlm)  [^4] https://huggingface.co/microsoft/wavlm-base-plus-sv model to extract speaker embeddings from both the prompt and synthesized speech, and compute the cosine similarity between them to assess speaker similarity. Finally, DNSMOS (dnsmos) is applied in TTS and VC to assess speech naturalness. It is a no-reference neural metric that predicts mean opinion scores (MOS) on a 5-point Likert scale, approximating human judgments of audio quality.
Subjective Evaluation Metrics
Considering the cost of human evaluation, we perform subjective assessments primarily for the TTS and VC tasks. For each task, assessments are performed from two perspectives: speech naturalness/quality (QMOS) and speaker similarity (SMOS). (1)QMOS evaluates the speech’s overall quality, naturalness, and clarity, while disregarding differences in speaking voice and speaking style such as emotion and prosody. (2)SMOS assesses how closely the speaker identity (i.e., timbre) in the generated speech matches the that in the reference speech, ignoring content, grammar, or audio fidelity of the generated speech. The scoring follows a 5-point Likert scale (5: excellent, 4: good, 3: fair, 2: poor, 1: bad). For each task, 15 audio samples are randomly selected and rated by 20 human evaluators.
DualSpeechLM Configuration
Model Configuration
Embedding dim in AcousticGPT 1024
Number of heads in AcousticGPT 8
Max context length (#tokens) 2000
LoRA rank 16
Optimization Configuration
Global batch size 72
Optimizer AdamW
Optimizer hyperparameter $\beta_{1}=0.9,\beta_{2}=0.95$
Warmup Steps 3500
Peak learning rate 1e-4
Minimum learning rate 1e-5
Learning rate decay 5e-2
Numerical precision bf16
Table 11: DualSpeechLM Configuration.
You are a professional evaluator for Question Answering (QA) tasks.
Firstly, you will be given the following content:
- speech content (from which questions will be asked based on the speech content),
- questions,
- reference answers,
- and then I will give you some candidate answers generated by different models sequentially.
You need to evaluate the overall quality of every models output across all samples.
Please assess the models answers based on the following criteria:
(1) Relevance (1-100): How well do the answers address the questions?
(2) Fluency (1-100): How natural and grammatically correct are the answers?
(3) Correctness (1-100): How accurately do the answers reflect the information in the reference answers?
Compare and evaluate these candidate answers generated by different models.
After reading through all the sample questions, please give a final score (1-100
points) for each candidates answers, taking all three criteria into account. You
should also write a brief explanation for each score.
Figure 6: Prompt for QA Evaluation using ChatGPTScore.
Appendix E D. Model Configuration
USTokenizer Configuration
[…]
DualSpeechLM Configuration
To verify the generalization of USToken and the effectiveness of the DualSpeechLM framework on different backbones, we conduct experiments using text LLMs of Vicuna-7B [^6] https://huggingface.co/lmsys/vicuna-7b-v1.1 and Phi-3B [^7] https://huggingface.co/microsoft/Phi-3.5-mini-instruct (abdin2024phi3technicalreporthighly). For AcousticGPT, we adopt 6 multi-head self-attention layers, each featuring an embedding dimension of 1024 and 8 attention heads. Causal masking is applied in each layer to effectively model sequential dependencies. During training, we set the weighting hyperparameters both $\lambda$ and $\xi$ in $\mathcal{L}_{\text{generation}}$ to 1. All DualSpeechLM models are trained on 4 NVIDIA A100-40GB GPUs for 60,000 steps. The detailed model and optimization configurations are summarized in Table 11. Furthermore, it is important to note that in SpeechGPT (zhang2023speechgpt) (Table 1 and Table 2), certain tasks, such as S2TT and T2ST, are not inherently supported. During inference, the S2TT task is accomplished by first applying ASR to transcribe speech into text, followed by translating the resulting English text into German. For the T2ST task, we first translate the Spanish and French text inputs into English, and subsequently perform TTS to synthesize speech.
Appendix F E. Semantic Comparison of USTokenizer
Task Prompt
ASR Recognize the speech and give me the transcription.
TTS Please read this sentence out loud.
VC Without altering the spoken content, transform the speaker’s voice in this speech to match the target voice.
T2ST Please translate the [sourcelanguage][source\;language] text into [targetlanguage][target\;language] speech.
SC Please listen to the speech content and provide a spoken answer to the question.
SQA Based on the content, provide a text-based answer to the question.
S2TT Please translate the [sourcelanguage][source\;language] speech into [targetlanguage][target\;language] text transcription.
SER Please describe the emotion of the speaker.
Table 12: Task prompts used in multi-task evaluation.
Evaluation with Recent Tokenizers
To further validate the semantic capability of our USTokenizer, we retrain automatic speech recognition (ASR) models using several recent tokenizers on Librispeech, including TAAE (parkerscaling), MimiCodec (defossez2024moshi), and CosyVoice-2 (du2024cosyvoice2), under identical experimental setups. For fairness, we only use the first quantization layer of MimiCodec, which is distilled from the first layer of WavLM (chen2022wavlm). As shown in Table 13, USTokenizer achieves the lowest Word Error Rate (WER) on both clean and noisy speech conditions. These results demonstrate that USTokenizer provides a more effective abstraction of high-level information, indicating its stronger ability in semantic preservation.
Model Clean Other
$wer\downarrow$
TAAE 13.24 29.49
MimiCodec 5.96 14.54
CosyVoice2 6.29 15.19
USTokenizer(Ours) 4.22 9.71
Table 13: The Comparison with Recent Tokenizers.
Comparison with Baichuan-Audio Tokenizer
We also compare USTokenizer with Baichuan-Audio (li2025baichuanaudio), a recent unified framework for end-to-end speech interaction. Baichuan-Audio adopts an 8-layer Residual Vector Quantization (RVQ) scheme optimized for Mel-spectrogram reconstruction. In contrast, USTokenizer employs a single-layer VQ optimized for self-supervised representation reconstruction, which is simpler, more LLM-friendly, and emphasizes semantic preservation. As summarized in Table 14, Baichuan-Audio exhibits a higher WER, possibly because Mel reconstruction prioritizes acoustic fidelity, which in turn diminishes semantic retention to some extent. By contrast, USTokenizer reconstructs self-supervised representations, achieving a lower bitrate and richer semantic encoding.
Model Frame Rate Token Bitrate Clean/Other
HzHz $num/s$ kbpskbps $wer\downarrow$
Baichuan-Audio 12.5 100 1.08 4.53/10.65
USTokenizer(Ours) 25 25 0.25 4.22/9.71
Table 14: The Comparison with Baichuan-Audio on Librispeech.
Appendix G F. Computational Overhead Analysis
To quantify the additional computational cost introduced by USTokenizer and DualSpeechLM, we report the relative changes in memory and training time under different configurations. As shown in Table 15, the text-LLM module in USTokenizer introduces moderate overhead but remains manageable, as it is unused during inference and converges quickly when applied to DualSpeechLM. Furthermore, DualSpeechLM demonstrates higher training efficiency owing to the low-bitrate token design of USTokenizer, which reduces both memory and time consumption compared with baselines.
Model Setting Memory(%) Time(%)
USTokenizer vs w/o Text LLM $\uparrow$ 288 $\uparrow$ 76
DualSpeechLM vs Baseline-Semantic $\downarrow$ 26 $\downarrow$ 32
DualSpeechLM vs Baseline-Acoustic $\downarrow$ 17 $\downarrow$ 31
Table 15: Training overhead of USTokenizer and DualSpeechLM.
Overall, these results indicate that the additional text-LLM module in USTokenizer slightly increases training cost but does not affect inference efficiency, while DualSpeechLM achieves better computational efficiency.
Appendix H G. Prompts for different tasks
Table 12 summarizes the prompts used for different tasks in training and evaluation of our DualSpeechLM.
Appendix I H. Discussion
Although Figure 1 shows that understanding and generation tasks can mutually enhance each other in our framework, a closer look at subfigures (a) Right and (b) Left reveals an interesting asymmetry: increasing the data for understanding significantly boosts generation performance, whereas increasing generation data yields only limited improvements in understanding. This may be because the understanding task is inherently ‘stronger’ in our setting. When more understanding data is provided, the text LLM can better learn the mapping between text tokens and DualTokens, leading to more accurate DualToken representations. These improved semantic representations in turn enhance AcousticGPT’s predictions, thereby improving generation performance. Conversely, even when generation data increases, its relatively weaker semantic quality limits the benefit it can bring to the understanding task. Nevertheless, our semantic supervision loss mitigates this issue by explicitly improving the model’s semantic representation capability during generation. This enables generation data to still have a positive, albeit smaller, effect on understanding. In contrast, the baseline model in subfigure (b) Left relies solely on acoustic tokens, which lack sufficient semantic information, resulting in poor performance on understanding tasks.